OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 3 and a y-intercept of -2.

OpenStudy (kymber):

The equation is y = mx + b m is the slope and b is the y-intercept Plug in what you know and done!

OpenStudy (anonymous):

i dont know

OpenStudy (kymber):

You don't know what? Plug what you know into y = mx + b I just explained it!

OpenStudy (anonymous):

i dont know where the stuff im suppose plug in goes

OpenStudy (kymber):

Read! "m is the slope and b is the y-intercept" so y = mx + b y = (SLOPE)x + (Y-INT) Look in the problem for the slope and the y-intercept!

OpenStudy (anonymous):

no need to be rude but is the answer y=3x-2

OpenStudy (kymber):

I am sorry if I came off as rude, I am certainly not trying to be. That is correct.

OpenStudy (anonymous):

ty and its fine have a good day